OpenStudy (anonymous):

simplify completely i^99 (imaginary number)

OpenStudy (anonymous):

ok 1 question here. Do you know i = root of -1, i^2 = -1 and hence i^4 = 1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok if i^4 is 1. Is it true that i^ any multiple of 4 = 1 ?

OpenStudy (raden):

i^99 = i^98 * i = (i^2)^49 * i = ((sqrt(-1))^2)^49 * i = (-1)^49 * i = ...

OpenStudy (anonymous):

If the exponent is odd, it equals -1, If the exponent is even, it equals 1. so answer is -1

OpenStudy (anonymous):

@Luis_Rivera it would be -1 iota since Iota still remains.